OpenStudy (openstudyrocks5*):

Write an inequality for this sentence. Then solve the inequality. Fifteen less than a number is at most 6.

OpenStudy (openstudyrocks5*):

@iGreen. @silverr

OpenStudy (anonymous):

fifteen less than a # is at most 6. y - 15 < 6 think you can try from here?

OpenStudy (openstudyrocks5*):

y = 21?

OpenStudy (anonymous):

the sign is not = what do you think it is? did you divide? should it be flipped?

OpenStudy (anonymous):

@OpenStudyRocks5* ?

OpenStudy (openstudyrocks5*):

@silverr Thinking. Im not very good at inequality's. :(

OpenStudy (anonymous):

okay :) it is only flipped when you divide by a negative. you're not dividing at all in this one, and not doing anything with a negative. therefore ?

OpenStudy (openstudyrocks5*):

I feel like an idoit :|

OpenStudy (openstudyrocks5*):

@iGreen.

OpenStudy (anonymous):

no, your not ! you just need practice is all :)

OpenStudy (anonymous):

the sign stays the same because you are doing nothing.

OpenStudy (openstudyrocks5*):

@iGreen.

OpenStudy (anonymous):

The inequality is wrong.. I believe it's: \(y - 15 \le 6\) Our first step is to add 15 to both sides, what do you get?

OpenStudy (openstudyrocks5*):

Now when you say both sides what do you mean? 15 and 6? I haven't got that yet.

OpenStudy (anonymous):

Yeah, we are adding 6 + 15

OpenStudy (openstudyrocks5*):

6 + 15 = 21

OpenStudy (anonymous):

Yes, so we have: \(y \le 21\)

OpenStudy (openstudyrocks5*):

But the answers are longer than that so it cant be that... here ill give some of the answers.... @iGreen.

OpenStudy (openstudyrocks5*):

x + 15 ≤ 6; x ≤ 21. Which is probally right cause of what you just said (y ≤ 21)

OpenStudy (anonymous):

Yep, you got it..but the first equation is wrong..it's not: \(x + 15 \le 6\) It's: \(x - 15 \le 6\)

OpenStudy (openstudyrocks5*):

so x - 15 ≤ 6; x ≤ 21

OpenStudy (anonymous):

Yep, you got it.